OpenStudy (anonymous):

Please, i desperately need help. It's winter break, my teacher is not able to help because of vacation and no video or math tips that i know of is helping me right now. I really want to understand this!

jimthompson5910 (jim_thompson5910):

The roots are -1+4i and -1 - 4i so we know x = -1+4i or x = -1 - 4i ------------------------------------------------------- now add 1 to both sides for both equations x+1 = -1+4i+1 or x+1 = -1 - 4i+1 x+1 = 4i or x+1 = -4i ------------------------------------------------------- Then square both sides to get rid of the 'i' term x+1 = 4i or x+1 = -4i (x+1)^2 = (4i)^2 or (x+1)^2 = (-4i)^2 (x+1)^2 = 16i^2 or (x+1)^2 = 16i^2 (x+1)^2 = 16i^2 (x+1)^2 = 16(-1) (x+1)^2 = -16 ------------------------------------------------------- making sense so far?

OpenStudy (anonymous):

so @AlexandervonHumboldt2 I multiplied that and got 3x^2-15x. i must be doing something wrong :(

OpenStudy (anonymous):

@jim_thompson5910 WOW yes it does!

jimthompson5910 (jim_thompson5910):

I'm glad it does, let's keep going

jimthompson5910 (jim_thompson5910):

Now expand out the left side (x+1)^2 = -16 (x+1)(x+1) = -16 x(x+1)+1(x+1) = -16 x*x+x*1 + 1*x + 1*1 = -16 x^2 + x + x + 1 = -16 x^2 + 2x + 1 = -16 Then you add 16 to both sides x^2 + 2x + 1 = -16 x^2 + 2x + 1+16 = -16+16 x^2 + 2x + 17 = 0

jimthompson5910 (jim_thompson5910):

So, if you were to solve x^2 + 2x + 17 = 0 for x, you would get x = -1+4i or x = -1 - 4i I recommend you use the quadratic formula to check that claim.

OpenStudy (alexandervonhumboldt2):

k

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt (b ^(2)-4ac }{ 2a }\] that is the quadratic equation, correct? @jim_thompson5910 and thank you!

jimthompson5910 (jim_thompson5910):

if you meant to say \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] then yes, that is the quadratic formula

OpenStudy (anonymous):

sorry, yes! that's what i was trying to say. many thanks!

jimthompson5910 (jim_thompson5910):

no worries, you have to say b^{2} to mean \(\Large b^{2}\) if you aren't sure how to type something up in the equation format, and you see someone else write it in the equation format, you can right-click on the equation they wrote and go to "show math as" and then go to "TeX commands"

OpenStudy (jhannybean):

Another way to cross check is the completing the square method: \[x^2 +2x +17=0\]\[(x^2+2x)+17=0\]\[(x+1)^2 +17-1=0\]\[(x+1)^2=-16\]\[x+1 = \pm \sqrt{-16}\]\[x=-1\pm \sqrt{-1}\sqrt{16}\]\[x=-1\pm 4i ~~\checkmark \]

OpenStudy (anonymous):

i tried that tip, and it worked! i can't thank you enough :) @jim_thompson5910 @Jhannybean so that's how you check it, i see! thanks so much for the tips guys :) I can use these to find out how to work out my other problems

OpenStudy (jhannybean):

@jim_thompson5910 Let's try this by using Vieta's Formula! :)

OpenStudy (jhannybean):

So we can let \(\alpha = -1+4i\) and \(\beta = -1-4i\) ?

OpenStudy (jhannybean):

We are trying to get to: \(x^2 -(\alpha + \beta)x +\alpha\beta = 0\) So the sum: \(\alpha + \beta = -1+4i -1-4i = -2 ....\) stuck :\

jimthompson5910 (jim_thompson5910):

what is \(\Large \alpha*\beta\) equal to in this case?

OpenStudy (jhannybean):

\[\alpha\beta = (-1-4i)(-1+4i) = 1-4i+4i-16i^2 = 17\]

jimthompson5910 (jim_thompson5910):

shortcut: \[\Large (a+bi)*(a-bi) = a^2 + b^2\]

OpenStudy (jhannybean):

is it really??? oh my gosh.

jimthompson5910 (jim_thompson5910):

yes, \[\Large (a+bi)(a-bi) = a(a-bi)+bi(a-bi)\] \[\Large (a+bi)(a-bi) = a^2-abi + abi - b^2i^2\] \[\Large (a+bi)(a-bi) = a^2 +0abi - b^2(-1)\] \[\Large (a+bi)(a-bi) = a^2 + b^2\]

OpenStudy (jhannybean):

wait.. then \((a+b)(a-b) = a^2-b^2\) ?

jimthompson5910 (jim_thompson5910):

the 'i's change it a bit though

OpenStudy (jhannybean):

oh, because \(i^2 = -1\)

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (jhannybean):

nice! :D

OpenStudy (jhannybean):

So now I need to figure out the sum..\(\alpha +\beta\)

jimthompson5910 (jim_thompson5910):

so, \[\Large x^2 -(\alpha + \beta)x +\alpha\beta = 0\] \[\Large x^2 -(-2)x +17 = 0\] \[\Large x^2 + 2x +17 = 0\]

jimthompson5910 (jim_thompson5910):

you already figured out alpha + beta

OpenStudy (jhannybean):

Ohh... \(+4i -4i = 0\) Haha, i see.

OpenStudy (jhannybean):

I thought for some reason we needed to substitute \(i=\sqrt{-1}\) in there.

OpenStudy (jhannybean):

Thanks jim!!!!!! If i could give you another medal for this I def would.

jimthompson5910 (jim_thompson5910):

sure thing

ganeshie8 (ganeshie8):

\[(a+bi)(a-bi) = a^2 - (bi)^2 = \cdots \]

OpenStudy (jhannybean):

Oh that's another cool way to look at it!!

ganeshie8 (ganeshie8):

\[(\heartsuit + \clubsuit) (\heartsuit - \clubsuit) ~~ =~~\heartsuit^2 - \clubsuit^2 \]

Nnesha (nnesha):

\(\huge\color{green}{O.O~~~~ AWESOME :)}\)

Nnesha (nnesha):

\(\huge\color{green}{(O.O + *.*)(O.O - *.*)}\) \(\huge\color{red}{=}\) \(\huge\color{green}{O.O^2 - *.*^2}\) shorry ::D :D